# Dist Loss: Enhancing Regression in Few-Shot Region through Distribution Distance Constraint

### v1

@@ A.4.3 Performance of Dist Loss across different imbalanced ratios
We validated the effectiveness of Dist Loss by varying the imbalance ratios of the ECG-Ka-DIR dataset. The data distribution diagrams are shown in Figure 4, and the corresponding results in the few-shot regions are presented in Table 13. Across eight datasets with different imbalance ratios, our method achieved the best performance in six cases and the second-best performance in the remaining two. These results collectively demonstrate the robustness of our approach across varying levels of data imbalance.
Table 13: Performance of Dist Loss in the few-shot regions across eight datasets derived from the ECG-Ka-DIR dataset with varying imbalance ratios, with the best results highlighted in bold.
@@
Figure 4: Data distribution diagrams for the eight datasets derived from the ECG-Ka-DIR dataset with varying imbalance ratios.

### v2

@@ A.4.3 Performance of Dist Loss across different imbalanced ratios
We validated the effectiveness of Dist Loss by varying the imbalance ratios of the ECG-K-DIR dataset. The data distribution diagrams are shown in Figure 5, and the corresponding results in the few-shot regions are presented in Table 13. Across eight datasets with different imbalance ratios, our method achieved the best performance in six cases and the second-best performance in the remaining two. These results collectively demonstrate the robustness of our approach across varying levels of data imbalance.
Table 13: Performance of Dist Loss in the few-shot regions across eight datasets derived from the ECG-K-DIR dataset with varying imbalance ratios, with the best results highlighted in bold.
@@
Figure 5: Data distribution diagrams for the eight datasets derived from the ECG-K-DIR dataset with varying imbalance ratios.
@@ A.5 Performance of Dist Loss on the GM Metric
A.5.1 Performance visualization of Dist Loss
Figure 7 illustrates the performance comparison of three methods—vanilla model, LDS, and Dist Loss—on the IMDB-WIKI-DIR, AgeDB-DIR, and ECG-K-DIR datasets. As observed, Dist Loss achieves predictions that are systematically closer to the diagonal line ($y=x$, where the predicted values align with the ground truth values) across all datasets, indicating improved accuracy. The task on the ECG-K-DIR dataset, which involves estimating blood potassium concentration from single-lead ECG signals, is particularly challenging due to the inherently limited information provided by single-lead ECGs. This limitation exacerbates the regression dilution phenomenon, leading to larger deviations from the diagonal across all methods. Despite this difficulty, Dist Loss demonstrates superior performance, underscoring its robustness and effectiveness in addressing regression tasks with imbalanced and noisy data.
(a) Scatter plots on the IMDB-WIKI-DIR dataset: vanilla (left), LDS (middle), Dist Loss (right).
(b) Scatter plots on the AgeDB-DIR dataset: vanilla (left), LDS (middle), Dist Loss (right).
(c) Scatter plots on the ECG-K-DIR dataset: vanilla (left), LDS (middle), Dist Loss (right).
Figure 7: Performance visualization of vanilla, LDS, and Dist Loss